OpenStudy (anonymous):

the length of a rectangle is 2 cm less than 7 times the width. the perimeter is 60 cm. find the width and the length.

OpenStudy (chriss):

7w-2+w=30 8w-2=30 8w=32 w=4 width = 4 length = 7w-2 = 26

OpenStudy (anonymous):

thanks